OpenStudy (anonymous):

How do i factor x^2-225?

OpenStudy (anonymous):

Look for the square root of 225

OpenStudy (anonymous):

I posted it on your previous question

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

(x-1)(x+1)=x^2+x-x-1=x^2 -1 and yes it is 15... use this same concept

OpenStudy (anonymous):

√225=15

OpenStudy (anonymous):

so (x+15)(x-15)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

49x^2-25? Do i get two factors of -25 and equal 49 when added? or?

OpenStudy (anonymous):

that* instead of and

OpenStudy (anonymous):

here you could use the box method I drew on the other question you had of (x-2)(x+6)

OpenStudy (anonymous):

i dont like that method lol.

OpenStudy (anonymous):

lol, ok lets see what we can do

OpenStudy (anonymous):

just tell me what to do. dont give me the answer though

OpenStudy (precal):

 also, known as a difference of squares